OpenStudy (anonymous):

How do you solve the inequality x^2(x-3)(x+1)/(x-5)(x+7) > or = to 0.... using a number line.

OpenStudy (anonymous):

\(\dfrac{x^2(x-3)(x+1)}{(x-5)(x+7)}\geq0\) One step would be to figure out which values for x make this inequality undefined.

OpenStudy (anonymous):

sorry there is a network error but the first thing to do, you need to cross multiply

OpenStudy (anonymous):

cross multiply? I plugged points in on a number line (-7, -1, 0, 3 and 5) I just need to figure out which sections I need to shade in

OpenStudy (anonymous):

I will admit that I'm not sure how to solve this problem yet, but I don't see how cross multiplication will work.

OpenStudy (anonymous):

are you still there buddy, it seems everthing is okay now

OpenStudy (anonymous):

The original poster closed the question, but I am still curious about the answer....

OpenStudy (anonymous):

do you now how to get those values that poster got

OpenStudy (anonymous):

Yes, set each factor equal to 0.

OpenStudy (anonymous):

5 and -7 make the inequality undefined, so those should be open circles.

OpenStudy (anonymous):

so after that, you subtitude you values into the equation you got after cross multiplication

OpenStudy (anonymous):

I still don't understand how cross multiplication works in this case.

OpenStudy (anonymous):

, your number line will be like this

OpenStudy (anonymous):

sorry G, what do you mean you still dont understand how cross multiplication comes in this case

OpenStudy (anonymous):

hey K, are you there

OpenStudy (anonymous):

Cross multiplication usually means multiplying across the equal sign. In this case, that would end up with 0=0.

OpenStudy (anonymous):

no, if you cross multiply\[x^2(x-3)(x+1)=0\]

OpenStudy (anonymous):

Right, because the fraction would be \(\dfrac{0}{1}\) and not \(\dfrac{0}{0}\). Got it now. Thanks.

OpenStudy (anonymous):

good, i think now we can carry on

OpenStudy (anonymous):

K are you there